OpenStudy (anonymous):

For what value of the constant c is the function ƒ continuous on (-∞,∞) where ƒ(y)= {cy +3 if y (-∞, 3] {cy^2 -3 if (3,∞) ?

OpenStudy (dumbcow):

for f(y) to be continuous f(3) = c(3^2) -3 3c +3 = 9c -3 6c = 6 c = 1

OpenStudy (anonymous):

thanks!!